Blog from March, 2022

Teardown day 9

(late posting of this message from Friday)

On site: Steve, Chris, Tony, Clayton, Dexter (Matt left today)

Removed all stuff from DCS tower site, including spooling cables.  Removed all stuff from storage.  Final packing of base and both Penskes.  Base departed.  We departed – one Penske went to St. George on Friday night, the other arrived in Boulder.

CFACT is over!  (Though I still need to send follow-ups to landowners.)

Teardown day 8

Warm and some clouds.  Sebastian came to help and collect the remaining UU equipment.

CC: Used the forklift and lots of muscle to completely teardown this site.

UP: Removed this site.  Again, the soil heat flux plate was tilted off horizontal by about 5 deg and sensors in general seemed deeper, likely due to the organic layer (grasses) that makes the surface difficult to define (especially when frozen).

MH: Matt just tugged and both stakes came free.  Took less time than it took me to text Chris on another subject!

Organized/packed base and Penskes.  Still have a bit to do.

Remaining task: remove stuff (including transformers and job boxes) left at DCS and sort.  Then watch the base and maybe TTs go byebye.

Teardown day 7

Getting warm.  All of us were in shirts – no jackets or sweaters – and ran out of water.

Rented the forklift and used it to remove all fencing and posts (including some left by the PIs) at both supersites.  Pulled all equipment out of PRS, but still left a bit at DCS, primarily because it became too muddy to get a vehicle in.  Will have to revisit that on another morning.

Kept the forklift for another 1/2 day, since we haven't yet gone to CC.

A tiring, but productive, day.  Thanks, all!

Teardown day 6

ISS teardown is done.  Bill, Lou, and Liz left today.  For a couple of days, the ISFS crew is me, Clayton, Chris, Tony, Matt, Dexter.

Started by organizing into the 2 Penskes that will leave here and emptying out storage.  Returned 16' Penske since no longer needed.

Next, tore down LC.  Found that one manfrotto tripod foot was in the soil plot, perhaps 8cm from the Qsoil.  Not ideal, but might be okay with the data.  The qsoil probe came out with a crack in the lower prong, that we thought might have occurred on install.  Should check the Qsoil data.  Other sensors came out okay, albeit with a LOT of work.  Now, Sebastian's sonic is boxed up.

Next emptied stuff, oriented Chris and Tony to the base organization, and the other 4 went to lunch.

Next, tore down DC.  Soils were even harder, since the popsicle sticks themselves were buried.  Hot water trick helped, but still took perhaps 3 man-hours.  All probes recovered.  The ISS CS125 was loaded into the lidar trailer.

Next, tore down MW (in just 1 hr!).  Soils were easier, since top 2 cm were thawed.  However, managed to chop through the Gsoil sensor cable twice (itself, already repaired earlier).  Other 3 sensors were okay.

A long day, but good progress and a great team.

Tomorrow is forklift day.  Plan to hit PRS, then DCS for fenceposts, power, and other cleanup, then remove CC.  I hope this works!  I guess the worst case is we rent it for a second day.